In 1 year's time, Mrs Farmer will be twice as old as she was 10 years ago. How old is she?

Answer by nyc_function(2741)

age now = x
age future = x + 1
age past = x - 10

equation: x + 1 = 2(x - 10)

x + 1 = 2(x - 10)
x + 1 = 2x - 20

x - 2x = -1 - 20

-x = -21

x = -21/-1

x = 21

She is 21 years old.